# Fast Robust Kernel Regression through Sign Gradient Descent with Early Stopping

### v6

@@ 5 Kernel Regression with the $\ell_{\infty}$ and $\ell_{1}$ Norms
In contrast to KRR, unless the data is uncorrelated, no closed-form solutions exist for Equations 13 and 14. However, the problems are still convex and solutions can be obtained using the iterative optimization method proximal gradient descent (Rockafellar,, 1976), which, in contrast to standard gradient descent, can handle the discontinuities of the gradients of the $\ell_{\infty}$ and $\ell_{1}$ norms.
@@ 5.1 Regularization through Early Stopping
The regularization strength needs to be carefully selected so that the model can generalize well to new data. While a too large regularization results in poor performance due to a too simple model, a too small regularization might result in a too complex model that incorporates the noise in the training data, and thus generalizes poorly to new data (overfitting). This is sometimes referred to as the bias-variance tradeoff. In practice, the regularization strength is usually selected by some sort of cross-validation, where parts of the data are kept aside during training and then used for validation once the model is trained. By training multiple models, all with different regularization strengths, one can select the model that performs best on the validation data.
@@
However, applying the proximal gradient descent algorithm might be computationally heavy, especially when evaluating several different regularization strengths. In Section 4, we showed how the solution of KGF with early stopping is very similar to that of KRR, with a later stopping time, $t$, corresponding to a smaller regularization strength, $\lambda$. Thus, running KGD until convergence once, all different stopping times, or regularization strengths, between $t=0$ (corresponding to $\lambda=\infty$) and $t=t_{\max}$ (corresponding to $\lambda=0$) are obtained through this single execution of the algorithm. (For more detailed descriptions of the connections between explicit regularization and early stopping, see the works by e.g. Friedman and Popescu, (2004), Tibshirani, (2015) or Allerbo et al., (2023).)
If there are iterative optimization algorithms corresponding to the $\ell_{\infty}$ and $\ell_{1}$ penalties, instead of solving the problem using proximal gradient descent once for each value of $\lambda$, all different early stopping regularization strengths could be obtained by one single call of that algorithm.
@@
The similarities between $\ell_{1}$ regularization (lasso) and the iterative optimization method forward stagewise regression (also known as coordinate descent, or, for infinitesimal step size, coordinate flow) are well studied, for instance by Efron et al., (2004) and Hastie et al., (2007). Efron et al., (2004) show that under certain conditions, including for uncorrelated data, the solutions paths of lasso and coordinate flow coincide, but even when these conditions do not hold, the solution paths tend to be very similar, where smaller correlations lead to larger similarities. By comparing Equations 10 and 11, or Equations 6 and 10 by Allerbo et al., (2023) for standard linear regression, we note that also in the case of $\ell_{2}$ regularization and gradient flow, the solution paths exactly coincide for uncorrelated data.
@@
The topic of a gradient-based algorithm similar to $\ell_{\infty}$ regularization does not seem to be as well studied. However, in Proposition 4, we show that the solutions of $\ell_{\infty}$ regularization and sign gradient flow coincide for uncorrelated data. This is shown using the constrained form of regularized regression, which, due to Lagrangian duality, is equivalent to the penalized form.
@@
Before stating the proposition, we first review the three gradient-based algorithms and their corresponding infinitesimal step size versions. For a loss function, $L(\cdot)$, that quantifies the reconstruction error, we denote the gradient and its maximum component at time step $k$ as
@@
where the absolute value of the gradient is evaluated element-wise. Then, the update rules of gradient descent/flow, sign gradient descent/flow, and coordinate descent/flow are stated in Equation 15, where $\eta$ denotes the optimization step size, and where $\bm{I_{m}}(t)$ is a diagonal matrix that defines which which gradient component(s) that is updated at time $t$ (for details about $\bm{I_{m}}$, see Allerbo et al., (2023)).
@@
Remark 1: For coordinate descent (flow), in each iteration (at each time), only the coordinate(s) corresponding to the maximal absolute gradient value is updated.
Remark 2: The name “coordinate descent” is sometimes also used for other, related, algorithms.
Proposition 4.
@@
    $\frac{d\bm{\beta}(t)}{dt}=-\text{sign}\left(\frac{\partial}{\partial\bm{\beta}(t)}\left(\left\|\bm{y}-\bm{X}\bm{\beta}(t)\right\|_{2}^{2}\right)\right),\ \bm{\beta}(0)=\bm{0}.$
@@
    $\frac{d\bm{\alpha}(t)}{dt}=-\text{sign}\left(\frac{\partial}{\partial\bm{\alpha}(t)}\left(\left\|\bm{y}-\bm{K}\bm{\alpha}(t)\right\|_{\bm{K}^{-1}}^{2}\right)\right),\ \bm{\alpha}(0)=\bm{0}.$
@@
Remark 1: For infinitesimal step size, sign gradient descent, Adam (Kingma and Ba,, 2014), and RMSProp (Hinton et al.,, 2012) coincide, up to constants of order $10^{-8}$ added for numerical stability (Balles and Hennig,, 2018). Thus, in the uncorrelated case, $\ell_{\infty}$ regularization also coincides with Adam and RMSProp with infinitesimal step size.
Remark 2: Similarly to for ridge regression and gradient flow, and for lasso and coordinate flow, the exact coincidence of $\ell_{\infty}$ regularization and sign gradient flow only holds for uncorrelated data. However, due to continuity, unless the correlations are too large, the solution paths still tend to be very similar.
@@
According to Proposition 5, KSGD and KCD correspond to feature space regression with the $\ell_{1}$ and $\ell_{\infty}$ loss functions, respectively. (Note that while we previously discussed different norms for the penalty term, Proposition 5 considers different norms for loss function.) Regression with the $\ell_{1}$ loss function, which is known as least absolute deviations, is a commonly used method for robust regression. Vice versa, regression with the $\ell_{\infty}$ loss function minimizes the maximum residual, i.e. more extreme observations contribute more to the solution than less extreme observations do.
@@ 6.1 Demonstrations on Synthetic Data
[…]
As proposed by Proposition 1, the KGF and KRR solution differs most where the non-regularized solution is large, and as suggested by Proposition 2, the KGF solution tends to lie closer to the observations than the KRR solution does.
[…]
@@ 6.2 Robust Kernel Regression on Real Data
Fast quantile regression, (Zheng,, 2022) implementation KQR-Z
Fast quantile regression, Tang et al., (2024) implementation KQR-T
In Tables 4 and 5, we present the computation time and $R^{2}$ on test data for the experiments. ($R^{2}:=1-\|\bm{y}-\bm{\hat{y}}\|_{2}^{2}/\|\bm{y}-\bar{y}\|_{2}^{2}\leq 1$, where $\bm{y}$ and $\bar{y}$ denotes the vector of observations and its mean, and $\bm{\hat{y}}$ denotes the vector of model predictions, measures the proportion of the variance in $\bm{y}$ that is explained by the model. $R^{2}=1$ corresponds to a perfect fit, while a model that predicts $\hat{y}_{i}=\bar{y}$ for all $i$ renders $R^{2}=0$; thus a negative value of $R^{2}$ is possible, but it corresponds to a model that performs worse than simply always predicting the mean). The difference between the two tables is that in Table 5, the presence of outliers is amplified by multiplying each element, $y_{i}$, in the response vector by $(1+|\varepsilon_{i}|)$, where $\varepsilon_{i}\sim\mathcal{C}(0,0.01)$ is Cauchy distributed. We note that amplifying the outliers only affects the relative performance between the robust and non-robust methods, not the relative performance among the robust methods.
@@
Comparing Tables 4 and 5, we note that KGD and KRR are much more sensitive to outliers than the robust methods are. We also note with amplified outliers, KRR tends to perform better than KGD in terms of test $R^{2}$, while in Table 4, the tendency is the opposite. This can probably be attributed to the fact that the KGF solution tends to lie closer to the observations than the KRR solution does, as suggested by Proposition 2. This is often beneficial, however not in the presence of outliers.
Table 3: Real data sets used for comparing the methods.
@@
Sound pressure of airfoils [^1] The data set is available at https://archive.ics.uci.edu/dataset/464/superconductivty+data. $1502\times 6$
@@
Energy consumption in steel production [^3] The data set is available at https://archive.ics.uci.edu/dataset/851/steel+industry+energy+consumption. $35040\times 7$
Critical temperature of superconductors [^4] The data set is available at https://archive.ics.uci.edu/dataset/464/superconductivty+data. $21263\times 82$
@@
Table 4: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers. The six robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 5: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers. The six robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@ A.1 Equivalent Formulations of Kernel Ridge Regression
In Lemma 4, we present three different objective functions, that all render the same solution for $[\bm{\hat{f}}^{\top},\bm{\hat{f}^{*}}{{}^{\top}}]^{\top}$ as in Equation 3. When expressing KRR in terms of $\bm{\alpha}\in\mathbb{R}^{n}$, it is enough to include $\bm{K}\in\mathbb{R}^{n\times n}$ in the expression to capture the dynamics of the kernel. However, when expressing KRR in terms of $[\bm{\hat{f}}^{\top},\bm{\hat{f}^{*}}{{}^{\top}}]^{\top}\in\mathbb{R}^{n+n^{*}}$, we need to introduce the larger kernel matrix
@@
Remark 1: We assert that $\bm{f^{*}}$ does not affect the reconstruction error by requiring that $\bm{\tilde{y}}-\bm{f^{*}}=\bm{0}$, We may, however, not define $\bm{\tilde{y}}$ to equal $\bm{f^{*}}$, since we do not want $\bm{\tilde{y}}$ to be considered when differentiating Equations 18b and 18c with respect to $\bm{f^{*}}$.
Remark 2: The term $\left\|\begin{bmatrix}\bm{f}\\ \bm{f^{*}}\end{bmatrix}\right\|^{2}_{(\bm{K^{**}})^{-1}}$, might appear a bit peculiar, including the inverse of the extended kernel matrix (which never needs to be calculated due to cancellations). However, it is very closely related to the expression for the reproducing kernel Hilbert space norm, $\|f\|^{2}_{\mathcal{H}_{k}}$, obtained when expressing the functions $f$ and $k$ in terms of the orthogonal basis given by Mercer’s Theorem: $f(\bm{x})=\sum_{i=1}^{\infty}f_{i}\phi_{i}(\bm{x})$ and $k(\bm{x},\bm{x^{\prime}})=\sum_{i=1}^{\infty}k_{i}\phi_{i}(\bm{x})\phi_{i}(\bm{x^{\prime}})$, with $\|f\|^{2}_{\mathcal{H}_{k}}:=\sum_{i=1}^{\infty}\frac{f_{i}^{2}}{k_{i}}$. Defining the vector $\bm{\tilde{f}}$ and the diagonal matrix $\bm{\tilde{K}}$ according to $\bm{\tilde{f}}_{i}=f_{i}$ and $\bm{\tilde{K}}_{ii}=k_{i}$, we obtain $\|f\|^{2}_{\mathcal{H}_{k}}=\sum_{i=1}^{\infty}\frac{f_{i}^{2}}{k_{i}}=\bm{\tilde{f}}^{\top}\bm{\tilde{K}}^{-1}\bm{\tilde{f}}=\|\bm{\tilde{f}}\|_{\bm{\tilde{K}}^{-1}}^{2}$.

### v7

@@ 5 Kernel Regression with the $\ell_{\infty}$ and $\ell_{1}$ Norms
In contrast to KRR, unless the data is uncorrelated, no closed-form solutions exist for Equations 12 and 13. However, the problems are still convex (strictly convex if $\bm{K}$ is strictly positive definite, which is the case for all kernels with infinite-dimensional feature expansions) and solutions can be obtained using the iterative optimization method proximal gradient descent (Rockafellar,, 1976), which, in contrast to standard gradient descent, can handle the discontinuities of the gradients of the $\ell_{\infty}$ and $\ell_{1}$ norms.
@@ 5.1 Regularization through Early Stopping
The regularization strength must be carefully selected for the model to generalize well on new data. While a too large regularization results in poor performance due to a too simple model, a too small regularization might result in a too complex model that incorporates the noise in the training data, and thus generalizes poorly to new data (overfitting). This is sometimes referred to as the bias-variance tradeoff. In practice, the regularization strength is usually selected by some sort of cross-validation, where parts of the data are kept aside during training and then used for validation once the model is trained. By training multiple models, all with different regularization strengths, one can select the model that performs best on the validation data.
@@
However, applying the proximal gradient descent algorithm might be computationally heavy, especially when evaluating several different regularization strengths. For some classes of models, there exist computationally efficient versions of leave-one-out and generalized cross-validation, but, as demonstrated in Appendix B, these do not perform well in this setting. Instead, we use another approach: In Section 4, we showed how the solution of KGF with early stopping is very similar to that of KRR, with a later stopping time, $t$, corresponding to a smaller regularization strength, $\lambda$. Thus, running KGD until convergence once, all different stopping times, or regularization strengths, between $t=0$ (corresponding to $\lambda=\infty$) and $t=t_{\max}$ (corresponding to $\lambda=0$) are obtained through this single execution of the algorithm. (For more detailed descriptions of the connections between explicit regularization and early stopping, see the works by e.g. Friedman and Popescu, (2004), Tibshirani, (2015) or Allerbo et al., (2023).) Analogously, if we can find iterative optimization algorithms that correspond to the $\ell_{\infty}$ and $\ell_{1}$ penalties, instead of solving the problem using proximal gradient descent once for each value of $\lambda$, all different early stopping regularization strengths could be obtained by one single call of that algorithm.
@@
The similarities between $\ell_{1}$ regularization (lasso) and the iterative optimization method forward stagewise regression (also known as coordinate descent, or, for infinitesimal step size, coordinate flow) are well studied, for instance by Efron et al., (2004) and Hastie et al., (2007). Efron et al., (2004) show that under certain conditions, including for uncorrelated data, the solutions paths of lasso and coordinate flow coincide, but even when these conditions do not hold, the solution paths tend to be very similar, where smaller correlations lead to larger similarities. Similarly, by comparing Equations 10 and 11, or Equations 6 and 10 by Allerbo et al., (2023) for standard linear regression, we note that also for $\ell_{2}$ regularization and gradient flow, the solution paths exactly coincide for uncorrelated data.
@@
The topic of a gradient-based algorithm similar to $\ell_{\infty}$ regularization does not seem to be as well studied. However, in Proposition 5.1, we show that the solutions of $\ell_{\infty}$ regularization and sign gradient flow coincide for uncorrelated data. This is shown using the constrained form of regularized regression, which, due to Lagrangian duality, is equivalent to the penalized form.
@@
Before stating the proposition, we first review the three gradient-based algorithms and their corresponding infinitesimal step-size versions. For a loss function, $L(\cdot)$, that quantifies the reconstruction error, we denote the gradient and its maximum component at time step $k$ as
@@
where the absolute value of the gradient is evaluated element-wise. The update rules of gradient descent/flow, sign gradient descent/flow, and coordinate descent/flow are stated in Equation 14, where $\eta$ denotes the optimization step size, and where $\bm{I_{m}}(t)$ is a diagonal matrix that defines which gradient component(s) that is updated at time $t$ (for details about $\bm{I_{m}}$, see Allerbo et al., (2023)).
@@
Remark 5.1.
For coordinate descent (flow), in each iteration (at each time), only the coordinate(s) corresponding to the maximal absolute gradient value is updated.
Remark 5.2.
The name “coordinate descent” is sometimes also used for other, related, algorithms.
Proposition 5.1.
@@
    $\frac{d\bm{\hat{\beta}}(t)}{dt}=-\text{sign}\left(\frac{\partial}{\partial\bm{\hat{\beta}}(t)}\left(\left\|\bm{y}-\bm{X}\bm{\hat{\beta}}(t)\right\|_{2}^{2}\right)\right),\ \bm{\hat{\beta}}(0)=\bm{0}.$
@@
    $\frac{d\bm{\hat{\alpha}}(t)}{dt}=-\text{sign}\left(\frac{\partial}{\partial\bm{\hat{\alpha}}(t)}\left(\left\|\bm{y}-\bm{K}\bm{\hat{\alpha}}(t)\right\|_{\bm{K}^{-1}}^{2}\right)\right),\ \bm{\hat{\alpha}}(0)=\bm{0}.$
@@
Remark 5.3.
For infinitesimal step size, sign gradient descent, Adam (Kingma and Ba,, 2014), and RMSProp (Hinton et al.,, 2012) coincide, up to constants of order $10^{-8}$ added for numerical stability (Balles and Hennig,, 2018). Thus, in the uncorrelated case, $\ell_{\infty}$ regularization also coincides with Adam and RMSProp with infinitesimal step size.
Remark 5.4.
Similarly to ridge regression and gradient flow, and lasso and coordinate flow, the exact coincidence of $\ell_{\infty}$ regularization and sign gradient flow only holds for uncorrelated data. However, due to continuity, unless the correlations are very large, the solution paths still tend to be very similar.
5.2 Properties of the Early Stopping Algorithms
@@
According to Proposition 5.2, KSGD and KCD correspond to feature space regression with the $\ell_{1}$ and $\ell_{\infty}$ loss functions, respectively. (Note that while we previously discussed different norms for the penalty term, Proposition 5.2 considers different norms for loss function.) Regression with the $\ell_{1}$ loss function, which is known as least absolute deviations, is a commonly used method for robust regression. Vice versa, regression with the $\ell_{\infty}$ loss function minimizes the maximum residual, i.e. more extreme observations contribute more to the solution than less extreme observations do.
Even for convex problems, a too large optimization step size will cause the solution to diverge. In Proposition 5.3, we present upper bounds for the step sizes, which guarantee that the solutions do not diverge.
Proposition 5.3.
When solving $\min_{\bm{\alpha}}\left\|\bm{y}-\bm{K}\bm{\alpha}\right\|_{\bm{K}^{-1}}^{2}$ for $\bm{\alpha}$ with
  • gradient descent, i.e. when $\bm{\hat{\alpha}}_{k+1}=\bm{\hat{\alpha}}_{k}+\eta\cdot\left(\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\right)$, then
    $\eta\leq\frac{2}{s_{\max}(\bm{K})}\implies\|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k+1}\|_{2}\leq\|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\|_{2},$
    where $s_{\max}(\bm{K})$ denotes the largest singular value of $\bm{K}$.
  • sign gradient descent, i.e. when $\bm{\hat{\alpha}}_{k+1}=\bm{\hat{\alpha}}_{k}+\eta\cdot\text{sign}\left(\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\right)$, then
    $\eta\leq\min_{1\leq i\leq n}\left(\left|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\right|_{i}\right)\implies\|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k+1}\|_{1}\leq\|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\|_{1},$
    where the absolute value is taken element-wise.
  • coordinate descent, i.e. when $\left(\bm{\hat{\alpha}}_{k+1}\right)_{m_{k}}=\left(\bm{\hat{\alpha}}_{k}\right)_{m_{k}}+\eta\cdot\text{sign}\left(\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\right)_{m_{k}}$, then
    $\eta\leq\min_{1\leq i\leq n}\left(\left|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\right|_{i}\right)\implies\|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k+1}\|_{\infty}\leq\|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\|_{\infty},$
    where $m_{k}$ denotes the coordinate with the largest absolute value, and where the absolute value is taken element-wise.
Remark 5.5.
Note that depending on which algorithm is used, the convergence is obtained for different norms, similar to in Proposition 5.2.
[…]
As proposed by Proposition 4.1, the KGF and KRR solutions differ most where the non-regularized solution is large, and as suggested by Proposition 4.2, the KGF solution tends to lie closer to the observations than the KRR solution does.
[…]
@@ 6.2 Robust Kernel Regression on Real Data
Fast quantile regression implemented by Zheng, (2022) KQR-Z
Fast quantile regression implemented by Tang et al., (2024) KQR-T
[…]
@@
Comparing Tables 4 and 5, we note that KGD and KRR are much more sensitive to outliers than the robust methods are. We also note with amplified outliers, KRR tends to perform better than KGD in terms of test $R^{2}$, while in Table 4, the tendency is the opposite. This can probably be attributed to the fact that the KGF solution tends to lie closer to the observations than the KRR solution does, as suggested by Proposition 4.2. This is often beneficial, however not in the presence of outliers.
Table 3: Real data sets, with corresponding reference when existing, used for comparing the methods.
@@
Sound pressure of airfoils [^2] The data set is available at https://archive.ics.uci.edu/dataset/291/airfoil+self+noise. $1502\times 6$
@@
Energy consumption in steel production (Sathishkumar et al.,, 2021) [^4] The data set is available at https://archive.ics.uci.edu/dataset/851/steel+industry+energy+consumption. $35040\times 7$
Critical temperature of superconductors (Hamidieh,, 2018) [^5] The data set is available at https://archive.ics.uci.edu/dataset/464/superconductivty+data. $21263\times 82$
@@
Table 4: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers. The six robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 5: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers. The six robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@ 7 Conclusions
Acknowledgments
The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
@@ A.1 Equivalent Formulations of Kernel Ridge Regression
In Proposition A.1, we present three different objective functions, that all render the same solution for $[\bm{\hat{f}}^{\top},\bm{\hat{f}^{*}}{{}^{\top}}]^{\top}$ as in Equation 3. When expressing KRR in terms of $\bm{\alpha}\in\mathbb{R}^{n}$, it is enough to include $\bm{K}\in\mathbb{R}^{n\times n}$ in the expression to capture the dynamics of the kernel. However, when expressing KRR in terms of $[\bm{\hat{f}}^{\top},\bm{\hat{f}^{*}}{{}^{\top}}]^{\top}\in\mathbb{R}^{n+n^{*}}$, we need to introduce the larger kernel matrix
@@
Remark A.1.
We assert that $\bm{f^{*}}$ does not affect the reconstruction error by requiring that $\bm{\tilde{y}}-\bm{f^{*}}=\bm{0}$, We may, however, not define $\bm{\tilde{y}}$ to equal $\bm{f^{*}}$, since we do not want $\bm{\tilde{y}}$ to be considered when differentiating Equations 17b and 17c with respect to $\bm{f^{*}}$.
Remark A.2.
The term $\left\|\begin{bmatrix}\bm{f}\\ \bm{f^{*}}\end{bmatrix}\right\|^{2}_{(\bm{K^{**}})^{-1}}$, might appear a bit peculiar, including the inverse of the extended kernel matrix (which never needs to be calculated due to cancellations). However, it is very closely related to the expression for the reproducing kernel Hilbert space norm, $\|f\|^{2}_{\mathcal{H}_{k}}$, obtained when expressing the functions $f$ and $k$ in terms of the orthogonal basis given by Mercer’s Theorem: $f(\bm{x})=\sum_{i=1}^{\infty}f_{i}\phi_{i}(\bm{x})$ and $k(\bm{x},\bm{x^{\prime}})=\sum_{i=1}^{\infty}k_{i}\phi_{i}(\bm{x})\phi_{i}(\bm{x^{\prime}})$, with $\|f\|^{2}_{\mathcal{H}_{k}}:=\sum_{i=1}^{\infty}\frac{f_{i}^{2}}{k_{i}}$. Defining the vector $\bm{\tilde{f}}$ and the diagonal matrix $\bm{\tilde{K}}$ according to $\bm{\tilde{f}}_{i}=f_{i}$ and $\bm{\tilde{K}}_{ii}=k_{i}$, we obtain $\|f\|^{2}_{\mathcal{H}_{k}}=\sum_{i=1}^{\infty}\frac{f_{i}^{2}}{k_{i}}=\bm{\tilde{f}}^{\top}\bm{\tilde{K}}^{-1}\bm{\tilde{f}}=\|\bm{\tilde{f}}\|_{\bm{\tilde{K}}^{-1}}^{2}$.